# NoiseSDF2NoiseSDF: Learning Clean Neural Fields from Noisy Supervision

### v1

@@ 6 Conclusion
In this work, we introduced NoiseSDF2NoiseSDF, a framework capable of recovering clean surfaces from noisy and sparse point clouds by leveraging paired noisy SDFs in a Noise2Noise denoising approach. We demonstrated that, at noise intensities of 0.01 and 0.02, the reconstructed surfaces produced by our method are notably cleaner and smoother, both quantitatively and visually, compared to previous works. In future research, we aim to explore additional applications of the NoiseSDF2NoiseSDF framework, such as scaling up point cloud sizes to enhance geometric detail recovery, or replacing components within the framework with alternative architectures trained from scratch, thus improving the model’s ability to represent point cloud noise and further enhancing denoising performance.
Appendix G Technical Appendices
Appendix G.1 Limitation
A limitation of NoiseSDF2NoiseSDF emerges when the noise level of input point cloud grows: surface completeness tends to degrade due to the denoising network’s over-smoothing, which erases fine-scale geometry. This shortcoming is compounded by a practical constraint in our pipeline: owing to limited compute, we rely on a pretrained point2SDF encoder that accepts at most 2048 points. The resulting low-resolution SDF representation caps the geometric fidelity that the decoder can ultimately recover, especially in highly corrugated regions. Future work will revisit both the smoothness prior and the encoder capacity to mitigate these effects without incurring prohibitive training cost.
Appendix G.2 Broader Impact
The proposed NoiseSDF2NoiseSDF framework enables denoising of signed distance fields directly from noisy supervision, making it applicable in settings where clean ground-truth data is difficult or costly to obtain. This can enhance the robustness of 3D reconstruction in domains such as robotics, augmented reality, cultural preservation, and medical imaging, where input data is often imperfect. By removing the dependency on clean supervision, the method also reduces dataset curation overhead and energy consumption. Nevertheless, the model may suppress subtle but meaningful geometric details or reflect biases present in the noisy training data, which warrants caution in critical applications.
@@ Appendix G.2 Broader Impact
Appendix G.3 Metrics Formula
@@
$\mathbf{F}_{1}(\tau)$ $=\frac{2\,\mathrm{precision}(\tau)\,\mathrm{recall}(\tau)}{\mathrm{precision}(\tau)+\mathrm{recall}(\tau)},\quad\tau=0.02,$
@@
Here $P$ and $Q$ are the sets of predicted and ground-truth 3D samples, with ||$\cdot$|| ⋅ | denoting cardinality. The Chamfer Distance ($\mathbf{CD}$) measures the average nearest-neighbor distance between the two sets in both directions. It is non-negative ($\geq 0$), and lower values indicate better geometric alignment.
@@
Appendix G.4 More Visualization
Figure Appendix G.1: More results on ABC, Famous

### v2

@@ 6 Conclusion
We introduced NoiseSDF2NoiseSDF, a framework that recovers clean surfaces from noisy, sparse point clouds in a Noise2Noise denoising approach. At noise levels of 0.01 and 0.02, our method produces cleaner and smoother surfaces, both quantitatively and visually, compared to previous works. In the future, we aim to explore additional applications of the NoiseSDF2NoiseSDF, such as scaling point cloud sizes for more complex geometry or replacing framework components with alternative architectures to improve noise representation and denoising performance.
7 Ethics statement
Our work, NoiseSDF2NoiseSDF, adheres to the ICLR Code of Ethics. The research does not involve human subjects or animals, and all datasets used—ShapeNet, ABC, Famous, and Real—are publicly available with proper citations. The models used in our experiments are also open-source and distributed under appropriate licenses. We confirm that there are no conflicts of interest related to this work. Additionally, we have ensured that the privacy and security of the data used in our experiments comply with relevant data protection guidelines.
8 Reproducibility statement
To ensure the reproducibility of our work, the code script is provided in the supplementary materials, see Appendix I.5. Furthermore, we will publicly release the complete source code repository along with step-by-step instructions to facilitate easy replication of our results.
For comprehensive details on the experimental setup—including the training procedure, evaluation metrics, test sets, computational resources, and hyperparameter configurations—readers are directed to Sections 4.2 and 5.
Appendix I Technical Appendices
Appendix I.1 Potential Impact and Applications
The proposed NoiseSDF2NoiseSDF framework enables denoising of signed distance fields directly from noisy supervision, making it applicable in settings where clean ground-truth data is difficult or costly to obtain. This can enhance the robustness of 3D reconstruction in domains such as robotics, augmented reality, cultural preservation, and medical imaging, where input data is often imperfect. By removing the dependency on clean supervision, the method also reduces dataset curation overhead and energy consumption.
@@ Appendix I.1 Potential Impact and Applications
Appendix I.2 The Use of Large Language Models
Our research does not involve large language models in any component of the core methodology. We only use LLMs to refine our writing, such as performing appropriate language translations, correcting grammatical errors, and helping adjust the phrasing of certain sentences in the paper.
Appendix I.3 Metrics Formula
@@
[…]
@@
Here $P$ and $Q$ are the sets of predicted and ground-truth 3D samples, with $|\cdot|$ denoting cardinality. The Chamfer Distance ($\mathbf{CD}$) measures the average nearest-neighbor distance between the two sets in both directions. It is non-negative ($\geq 0$), and lower values indicate better geometric alignment.
@@
Appendix I.4 More Visualization
Figure Appendix I.1: More results on ABC, Famous and Real under the noise level of 0.02
Appendix I.5 Other Supplementary Materials
To ensure the reproducibility of our work, we provide the code in a zip file, which includes scripts for model training, evaluation, and testing. Additionally, it contains the dataset source and the basic data preprocessing pipeline we used. For further details, please refer to the README.md file.